# Eigen Attention: Attention in Low-Rank Space for KV Cache Compression

### v1

@@ 5.3 Ablation Studies
Fine-tuning Steps: In Figure 4(b), we evaluate perplexity (PPL) for the C4 dataset evaluated on the MPT and Llama models for a range of fine-tuning steps. With just 500 fine-tuning steps, the PPL improves by 1.2 on average across all the models. 2000 steps are ideal, beyond which we observe an average increase of 0.61 in the perplexity score. We posit that this is due to model overfitting and reduced generalization ability.
@@ 6 Conclusion
In this work, we propose Eigen Attention, a novel technique that reduces the memory overhead associated with KV cache in LLMs. Eigen Attention is inspired by the observation that keys, queries, and values can be approximated using a few basis vectors, enabling the possibility of performing the attention operation in a low-rank space with minimal performance loss. To achieve this, we project the attention inputs into low-rank subspaces defined by a set of principal basis vectors computed offline using a calibration dataset in a one-shot manner. These projections are integrated into the weight matrices, which are utilized during inference to generate lower dimensional key and value matrices, thereby reducing the KV cache memory footprint. Our approach is orthogonal to existing KV cache compression strategies and can be used in synergy. Extensive experiments across a range of LLMs and language tasks demonstrate that Eigen Attention reduces the KV cache memory footprint by 40% with minimal performance loss and achieves up to a 60% improvement in attention operation latency compared to the standard attention baseline. Please refer to Appendix A.6 for important future directions for our work.
@@ Acknowledgements
This work was supported by the Center for the Co-Design of Cognitive Systems (COCOSYS), a DARPA sponsored JUMP center of Semiconductor Research Corporation (SRC), National Science Foundation and Intel.

### v2

@@ 5.3 Ablation Studies
Impact of calibration dataset: Table 7 shows the performance of MPT-7b model when different calibration datasets are used for obtaining low rank space of K, V vectors. We select samples from C4 Dodge et al. (2021) and Alpaca Taori et al. (2023) datasets for obtaining the low rank basis. We see that compressing using C4 calibration dataset achieves better perplexity on the C4 dataset itself while achieving slightly lower perplexity on Wikitext dataset (results in Table 2). Performance on zero shot tasks remains similar regardless of the calibration dataset choice.
@@
Fine-tuning Steps: In Figure 4(b), we evaluate perplexity (PPL) for the C4 dataset evaluated on the MPT and Llama models for a range of fine-tuning steps. With just 500 steps, the PPL improves by 1.2 on average across all the models. 2000 steps are ideal, beyond which we observe an average increase of 0.61 in the PPL. We posit that this is due to model overfitting on finetuning data.
@@ 6 Conclusion
In this work, we propose Eigen Attention, a novel technique that reduces the memory overhead associated with KV cache in LLMs. Eigen Attention is inspired by the observation that keys, queries, and values can be approximated using a few basis vectors, enabling the possibility of performing the attention operation in a low-rank space with minimal performance loss. To achieve this, we project the attention inputs into low-rank subspaces defined by a set of principal basis vectors computed offline using a calibration dataset in a one-shot manner. These projections are integrated into the weight matrices, which are utilized during inference to generate lower dimensional key and value matrices, thereby reducing the KV cache memory footprint. Our approach is orthogonal to existing KV cache compression strategies and can be used in synergy. Extensive experiments across a range of LLMs and language tasks demonstrate that Eigen Attention reduces the KV cache memory footprint by 40% with minimal performance loss and achieves up to a 60% improvement in attention operation latency compared to the standard attention baseline.
@@ Acknowledgements
This work was supported by the Center for the Co-Design of Cognitive Systems (COCOSYS), a DARPA sponsored JUMP center of Semiconductor Research Corporation (SRC), National Science Foundation and United States Department of Energy.